OpenStudy (anonymous):

The profits in hundreds of dollars, P(c), that a company can make from a product is modeled by a function of the price, c, they charge for the product: P(c) = –20c2 + 320c + 5,120. What is the maximum profit the company can make from the product?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

max at the vertex the first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case it \[\huge -\frac{320}{2\times (-20)}\]